OpenStudy (anonymous):

Help please D: A recursive function is shown below: f(1) = 5 and f(n) = f(n - 1) - 6; n > 1 Which of the following lists the terms in the sequence defined by this recursive function? Choices: 5, 1, 7, 13, 19, ... 5, -1, -7, -13, -19, ... 5, 11, 17, 23, 29, ... 5, -11, -17, -23, -29, ...

zepdrix (zepdrix):

Hmm, I'm pretty sure the kitty goes.... meow <.< but anyway, we have a recursive function for n values larger than 1. \(\Large\rm f(\color{orangered}{n})=f(\color{orangered}{n}-1)-6\) So for n=2 we have: \(\Large\rm f(\color{orangered}{2})=f(\color{orangered}{2}-1)-6\) \(\Large\rm f(\color{orangered}{2})=f(1)-6\) They told us that f(1)=5, yes?

zepdrix (zepdrix):

\(\Large\rm f(2)=5-6\) So we just plug it in. Understand how that works? :d

OpenStudy (anonymous):

sort of ._.

OpenStudy (anonymous):

not really lol, i'm confused

OpenStudy (anonymous):

can you explain to me how it works?

zepdrix (zepdrix):

Hmm function notation takes some getting used to :c

OpenStudy (anonymous):

oh

zepdrix (zepdrix):

See the way I color coded the n's in the first formula? You replace all of the n values at the same time. n corresponds to the term in the sequence. So n=1 corresponds to f(1) which gives us our first term 5. n=2 corresponds to f(2) which gives us our second term. \(\Large\rm f(n)=f(n-1)-6\) So we plug 2 in for our n's, \(\Large\rm f(2)=f(2-1)-6\) I feel like I'm repeating myself :c I can't think of another way to explain it.

OpenStudy (anonymous):

It's fine, I'll leave that question aside and do it later ^-^ thanks for helping though